# Translating Images to Road Network: A Sequence-to-Sequence Perspective

### v2

@@ V Conclusion
Appendix A RoadNet Sequence & Semi-Autoregressive RoadNet Sequence
Details of sequence construction The discretization of $v_{x},v_{y}$ is simply truncating the integer part. Integer representation of $v_{c}$ is Ancestor: 0, Lineal: 1, Offshoot: 2, Clone: 3. Discretizing $e_{px}$ and $e_{py}$ can be challenging since the Bezier control points may exceed the Bird’s Eye View (BEV) range, and their values may become negative. As a solution, we discretize $e_{px}$ and $e_{py}$ by applying the int function to $(e_{px}+10)$ and $(e_{py}+10)$, respectively, to avoid negative values. Figure LABEL:fig:seq_example shows a example of both RoadNet Sequence and Semi-Autoregressive RoadNet Sequence.
Appendix B Input and target sequence construction
Sequence embedding Each vertex-edge pair is represented by 6 integers. To prevent embedding conflicts between the 6 integers, we divide them into separate ranges which is shown in Table IX. As a default, we set the embedding size to $576$, which is sufficient to accommodate all the integer ranges.
TABLE IX: Embedding range of different integers.
Item Range
$v_{x},v_{y}$ $0\sim 199$
$v_{c}$ $200\sim 249$
$v_{d}$ $250\sim 349$
$e_{px},e_{py}$ $350\sim 569$
noise category $570$
EOS $571$
Start 575757572
n/a $573$
Synthetic noise objects technique The input sequence of RoadNet Sequence starts with a start token and the target sequence ends with an EOS token. The EOS token makes the model know where the sequence terminates, but the experiments have shown that it tends to cause the model to stop predicting early without getting the complete sequence. Inspired by [22], we use a similar sequence augmentation technique to alleviate the problem called the synthetic noise objects technique [22]. The technique composes of sequence augmentation and sequence noise padding. The sequence augmentation adds noise to the position of landmarks and the coefficient of centerlines in input sequence. Whereas, sequence noise padding is a padding technique. For input sequences, we generate synthetic noise vertices and append them at the end of the real vertices sequence. Each noise vertex includes random locations($v_{x}$, $v_{y}$), category($v_{c}$), index of parent($v_{d}$) and Bezier curve coefficient($e_{px}$, $e_{py}$). As for the target sequence, the EOS token is added to the end of the real vertices sequence. We set the target category($v_{c}$) of each noise vertex to a specific noise class(different from any of the ground-truth labels), and the remaining components($v_{x}$, $v_{y}$, $v_{d}$, $e_{px}$, $e_{py}$) of the noise vertex to the ”n/a” class, whose loss is not calculated in the back-propagation.
However, we only use sequence noise padding as sequence augmentation has been shown to cause a decrease in performance. The introduced modifications of the synthetic noise objects technique are illustrated in Figure 15
The padding rules of Semi-Autoregressive RoadNet Sequence are much the same as auto-regressive RoadNet Sequence. As mentioned in the main submission, we pad the 2-dimensional Semi-Autoregressive RoadNet Sequence to $[[y_{1,1},y_{1,2},\cdots,y_{1,L}],\cdots,[y_{M,1},y_{M,2},\cdots,y_{M,L}]]$, where $L$ is the maximum length of each sub-sequence and $M$ is the number of sub-sequences. The valid sub-sequences begin with a key-point. For each valid sub-sequence, we follow the same padding rules of RoadNet Sequence, except there isn’t a start token in an input sub-sequence because of the Key-point Prompt. We set the other sub-sequences to the ”n/a” class making the loss of these sub-sequences without a key-point not calculated.
[…]
Appendix C Precision-Recall curve

### v3

@@ V Conclusion and Discussions
Limitations. While our method demonstrates strong performance across various metrics and datasets, there are still some limitations that deserve future exploration. First, our current BEV Encoder relies on 2D projections, which may struggle to capture elevation or fine-grained vertical structures (e.g., tunnels, flyovers), especially in complex 3D scenes. Second, while SD-Maps offer valuable priors, their coarse resolution limits their effectiveness for precise landmark localization, which may restrict performance in high-precision navigation scenarios. We believe these challenges provide promising directions for future work, including exploring 3D aware representations, and leveraging richer map priors with higher spatial fidelity.